5. (This problem, due to Steven Strogatz, appeared in the New York Times.) Before going on vacation for a week, you ask your spacey friend to water your ailing plant. Without water, the plant has a 90% chance of dying. Even with proper watering, it has a 20% chance of dying. The probability that your friend will forget to water it is 20% (a) Whats the chance your plant will survive the week? (b) If your friend forgot to water it, whats the chance itll be dead when you return? (c) If its dead when you return, whats the chance that your friend forgot to water it?
Answer #1

a)P(plant survive)=P(watered and survive)+P(not watered and survive)

=(1-0.2)*(1-0.2)+0.2*(1-0.9)=0.66

b)

P(not survive|not watered)=0.90

c)

P(not watered|not survive)=P(not watered and not survive)/P(not survive)

=0.2*0.9/(1-0.66)=0.5294